the ten's digit of atwo digit number is three times the unit digit . the sum of the number and the unit digit is 32.find the number.

Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

Let the units digit be x
The tens digit will be 3x
x+3x = 32
4x = 32
x = 8

Answer by ikleyn(53765)

.
The ten's digit of a two digit number is three times the unit digit. The sum of the number and the unit digit is 32. Find the number.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let the units digit be x. 0 <= x <= 9.

Then tens digit will be 3x.
And the number itself will be 10*(3x)+x = 30x + x.

The fact that "the sum of the number and the unit digit is 32" means 

(30x + x) + x = 32,  or   32x = 32.

Hence, x = 1.  

And the number is 31.

Solved.